gimmie:

A ladder is placed 30 inches from a wall. It touches the wall at a height of 50 inches from the ground. The angle made by the ladder with the ground is degrees, and the length of the ladder is inches.

jhonyy9:

so like a first idea try draw this case

Luigi0210:

I think you're missing a few details

Narad:

The angle is \[\theta= \arctan (\frac{ 50 }{ 30 })\] The length of the ladder is \[\sqrt{50^{2}+30^{2}}\]